Solve the following systems: 

              2x -  y = 1
              6y - 2x = 9

Found 2 solutions by checkley71, AnlytcPhil:
Answer by checkley71(8403)

y=2x-1 then 6(2x-1)-2x=9 or 12x-6-2x=9 or 10x=15 or x=1.5 then y=2*2.5-1 or
y=5-1 or y=4

Answer by AnlytcPhil(1806)

Solve the following systems: 

              2x -  y = 1
              6y - 2x = 9

By substitution:

Solve the first for y:

               2x - y = 1
               2x - 1 = y
                    y = 2x - 1

Substitute (2x - 1) for y in the second:

              6y - 2x = 9
       6(2x - 1) - 2x = 9
        12x - 6  - 2x = 9
              10x - 6 = 9
                  10x = 15
                    x = 15/10
                    x = 3/2

Substitute 3/2 for x in y = 2x - 1
 
                    y = 2x - 1
                    y = 2(3/2) - 1
                    y = 3 - 1
                    y = 2

The solution is (x,y) = (3/2, 2)

Edwin
AnlytcPhil@aol.com